OpenStudy (anonymous):

Medal A physician orders enoxaparin for a patient with deep vein thrombosis. The in stock solution of enoxaparin is 300 mg / 3 mL and the dosage instructions are 1 mg/kg SC q.12.h., and the patient weighs 85kg. Provide complete, detailed instructions for proper drug administration

OpenStudy (lurker):

1 mg/kg a patient weighs 85 kg therefore he needs 85 mg

OpenStudy (lurker):

the solution you are giving him has 300 mg / 3 mL

OpenStudy (lurker):

how many mL do you have to give him for 85 mg then

OpenStudy (anonymous):

8.5 mL

OpenStudy (lurker):

no no, are you confused with anything i ve typed so far

OpenStudy (anonymous):

m not confused i just dont understand the 85 mg in mL how do you convert this

OpenStudy (lurker):

okays the solution you administer to the patient has 300 mg for every 3 mL

OpenStudy (lurker):

you have to give him x mL such that he only get 85 mg

OpenStudy (lurker):

you can look at the ratios there 300 mg for 3 mL means 100 mg for 1 mL

OpenStudy (lurker):

which is not low enough, can you think of how to get it to 85 mg for some other mL

OpenStudy (anonymous):

ok so we need to make 85mg to mL right

OpenStudy (lurker):

we arent turning mg to ml we are just trying to figure out the ratios here

OpenStudy (anonymous):

ohh i c

OpenStudy (lurker):

for example what we did in that middle step there how i got 100 mg for 1 ml is since 300/3 = 100/1 this is the same ratio in the same way 300/3=85/x what should x be to maintain the same ratio

OpenStudy (anonymous):

so we cross multiply to get it right

OpenStudy (lurker):

yep

OpenStudy (lurker):

solve for x

OpenStudy (lurker):